Can you help me with algebra 1 please I need help it I due tomorrow here is the problem
M+n=23/12
M-n=17/12

Answer by algebrahouse.com(1659)

m + n = 23/12
m - n = 17/12
-------------
2m = 40/12 {added the two equations}
m = 5/3 {divided each side by 2}

m + n = 23/12 {top original equation}
(5/3) + n = (23/12) {substituted (5/3), in for m, into top equation}
n = 1/4 {subtracted 5/3 from each side}

m = 5/3 and n = 1/4

For more help from me, visit: www.algebrahouse.com